# X-Fi: A Modality-Invariant Foundation Model for Multimodal Human Sensing

### v2

@@ 4.1 Experimental Setup
We train and evaluate our proposed X-Fi on two largest human sensing multimodal public datasets MM-Fi (Yang et al., 2024) and XRF55 (Wang et al., 2024), to assess its efficiency as a unified modality-invariant foundation model across diverse human sensing tasks, including Human Pose Estimation (HPE) and Human Activity Recognition (HAR). The details of the specific datasets utilized for each task are provided as follows. (i) HPE: MM-Fi consists of over 320k synchronized frames from 40 human subjects. Each synchronized frame consists of samples collected from 5 sensing modalities, including RGB images (I), Depth images (D), LiDAR point clouds (L), mmWave point clouds (R), and WiFi-CSI (W). The chosen annotation type is 3D whole-body keypoints, representing the positions of 17 human joints. (ii) HAR: MM-Fi and XRF55 are utilized for HAR task evaluation. MM-Fi consists of 27 action categories, including 14 daily activities and 13 rehabilitation exercises. We exclude WiFi-CSI modality from MM-Fi when addressing the HAR task, as each WiFi data sample spans only 100ms, which is insufficient to effectively capture the distinct data patterns associated with different actions. XRF55 includes 42.9K radio frequency samples collected by 3 sensing modalities, including mmWave Range-Doppler & Range-Angle Heatmaps (R), WiFi-CSI (W), and RFID phase series data (RF). XRF55 comprises 55 action classes collected from 39 human subjects, covering categories such as fitness activities and human-computer interactions.
@@
Different loss functions are applied as the training objectives for different human sensing tasks. For HPE, Mean-Squared-Error is chosen as the loss function, which measures the average squared euclidean distance between each of the 17 predicted whole-body key-points and their corresponding ground truth points in 3D space. For HAR, Cross-Entropy loss is applied between the predicted probability distribution and the ground truth distribution.
MM-Fi MPJPE PA-MPJPE
@@
To evaluate the performance of our proposed unified modality-invariant foundation model across various human sensing tasks and modality combinations, we formulate baseline models by adopting a feature fusion approach (Das et al., 2021; Marcard et al., 2016) and a decision-level fusion approach (Yang et al., 2024). For the feature fusion approach, first, we leverage the pre-trained modality-specific feature extractors in X-Fi to obtain features from each modality. Subsequently, for single-modal inputs, a MLP structure is applied to project the single-modal features into the task representation space. For multi-modal inputs, we concatenate the features from each modality to create a multi-modal feature, which is then processed by a MLP head with an average pooling layer to fuse and map the multi-modal features into the unified task representation space. We train a baseline model for each modality combination. And for different tasks, we optimize the parameters of each baseline model (including the pre-trained feature extractor) using different learning objectives. For the decision-level fusion approach, we first obtain the final-layer output of each individual modality. Then, we calculate the average of the outputs from all modalities in the combination to generate the fused output. The baseline results are presented in the Table 12.
@@
We conduct experiments on all humans sensing tasks and datasets with the following model settings. To standardize feature representations obtained by various modality-specific feature extractors, we apply linear projection units to map each modality feature representation to 32 features, each with feature dimension of 512. The positional embedding, derived from LiDAR 3D point cloud raw data due to its inherent capture of human body spatial information, is only provided in the multi-modal embedding if the LiDAR modality is present in the input. The backbone for both the cross-modal transformer and each modality-specific cross-attention transformer consists of a 1-layer decoder-only transformer structure with 8 multi-head attention heads and a scaling factor of 0.125. Each modality-specific key-value pair and the cross-modal embedding in X-Fusion block comprise 32 features with feature dimension of 512. The AdamW optimizer, with an initial learning rate of $1\times e^{-3}$ for HPE and $1\times e^{-4}$ for HAR, is chosen for model optimization. The training process is performed with a batch size of 16 on an NVIDA GeForce RTX 4090 GPU.
@@ 4.2 Quantitative Results
We compare our approach with the baseline methods regarding the performance on downstream human sensing tasks, including HPE and HAR. For convenience, we refer to the decision-level fusion baseline as “baseline1”, and the feature-level fusion baseline as “baseline2”. The results for X-Fi, as shown in Tables 12, represent average values obtained from multiple experiments. These values are derived by validating the performance of a single X-Fi model across different modality combinations that is trained using the proposed modality-invariant training strategy.
Human Pose Estimation.
We assess the performance of X-Fi on the MM-Fi dataset (Yang et al., 2024) by adopting two commonly-used metrics in HPE: Mean Per Joint Position Error (MPJPE) and Procrustes Analysis MPJPE (PA-MPJPE). For single-modal inputs, X-Fi improves the averaged MPJPE and PA-MPJPE by $6.0\%$ and $5.2\%$, respectively. For the dual-modality inputs, the averaged MPJPE and PA-MPJPE are improved respectively by $14.5\%$ and $14.8\%$ on baseline1, $23.7\%$ and $18.9\%$ on baseline2. Notably, for LiDAR and WiFi modalities, when used as individual inputs, X-Fi does not achieve good performance, with MPJPE difference of $-3.5\%$ and $+0.7\%$, respectively. However, when the two modalities are combined as inputs, the performance improves significantly, surpassing the baseline1 by $4.5\%$ and the baseline2 by $22.6\%$. For the combined input of three modalities, X-Fi demonstrates an average improvement of $17.0\%$ and $28.8\%$ in MPJPE, $19.5\%$ and $25.9\%$ in PA-MPJPE. For the combined input of four modalities, the results indicate an average enhancement of $17.0\%$ and $35.5\%$ in MPJPE, $21.6\%$ and $30.8\%$ in PA-MPJPE. For all five modalities combined input, MPJPE increases by $18.8\%$ and $35.1\%$, PA-MPJPE increases by $23.8\%$ and $34.3\%$.
Table 2: HAR performance on the MM-Fi and XRF55 datasets.
@@
R 82.1% 83.9% 1.9%
W 77.8% 55.7% -22.1%
RF 42.2% 42.5% 0.3%
R+W 86.8% 88.2% 1.4%
R+RF 71.4% 86.5% 15.1%
W+RF 55.6% 58.1% 2.6%
R+W+RF 70.6% 89.8% 19.2%
@@
The performance of X-Fi on the HAR task is evaluated by classification accuracy on both the MM-Fi (Yang et al., 2024) and XRF55 (Wang et al., 2024) datasets. On the MM-Fi dataset, X-Fi performs better than the baseline when handling multi-modal inputs, yielding an average improvement of 2.8%. But when dealing with single-modality inputs, its performance declines, dropping by 2.8%. On the XRF55 dataset, X-Fi only encounters challenges with WiFi-CSI single-modality input. It outperforms the baseline on all other modality combinations, achieving an average improvement of 6.7%. Notably, while X-Fi’s performance is suboptimal when using WiFi as a single input, when combined with other modalities, WiFi not only avoids hindering performance but also introduces unique information that enhances classification accuracy. For example, when X-Fi processes mmWave radar and RFID multimodal input, it outperforms the baseline by 15.1%. With the inclusion of WiFi, this improvement increases to 19.2%.
@@ 4.3 Analytics
To assess the impact of adjusting the independent existing probability of each modality in the modality existence list, we conduct experiments with different probability combinations on the XRF55 dataset for the HAR task. The XRF55 modality existence list contains the independent existence probabilities of mmWave Radar, WiFi-CSI, and RFID modalities in sequential order. As shown in Table 3, X-Fi encounters difficulties in dealing with WiFi-CSI. However, through gradually increase the exist probability of WiFi-CSI, the model is prioritised to fit WiFi-CSI modality during training, which yields significant improvement from $29.1\%$ to $55.7\%$ in handling single WiFi-CSI input. Additionally, we are surprised to find that prioritizing a particular modality does not degrade the model’s ability to extract features from the other two modalities. Consequently, the multimodal fusion results are also improved a lot.
@@
Method Transformer X-Fusion Transformer X-Fusion
@@
Table 4: Comparison between transformer-based fusion and X-Fusion on the HPE task..
Comparison between transformer-only fusion and X-Fusion.
To assess whether the X-Fusion module’s feature injection mechanism preserves distinctive features from each modality and enhances multi-modal fusion, we replaces X-Fusion in X-Fi with a basic transformer encoder consists of four stacked blocks. We conduct experiments on these variants, with a subset of results analyzed in Table 4 and the full findings provided in the appendix 5. As shown in Table 4, X-Fusion yields better performance on most of the modality combination input. First, when handling single modality input, transformer-based fusion struggles with challenging modalities, whereas X-Fusion effectively captures the distinctive features of each modality. For example, performance with LiDAR input improves by $21.3\%$. Next, in contrast to transformer-based fusion methods, X-Fusion leverages each modality’s strengths, yielding better fusion results than single-modality inputs. X-Fusion demonstrates superior performance compared to either the LiDAR or WiFi-CSI modality when used individually. Lastly, in the case of multi-modality inputs, the performance of transformer-based fusion is highly dependent on the presence of dominant modalities. For example, adding the RGB modality improves the performance of the LiDAR and WiFi-CSI multi-modal inputs by 119.7 mm with the transformer backbone, whereas the corresponding improvement with X-Fusion is 66.6 mm.
@@
To evaluate whether the proposed X-Fusion structure, which iterates multiple loops on the same X-Fusion block, is optimal for preserving and fusing multi-modal features, we design two variants. The first employs a layered architecture with four sequentially stacked X-Fusion blocks, where each layer learns new key-value pairs from the multimodal embedding of the previous layer. The second variant is similar, but uses the same key-value pairs across layers, learned from the embedding generated by the modality feature encoder. We conduct experiments on both variants, with full results provided in the appendix 5. On average, the first variant demonstrates a performance improvement of +1.27 mm for MPJPE and -0.7 mm for PA-MPJPE across all modality combinations. The second variant achieves an improvement of +0.93 mm for MPJPE and +0.03 mm for PA-MPJPE. Therefore, considering the computational efficiency and the minor performance differences, we retain the X-Fusion structure where an X-Fusion block is applied iteratively.
@@ 4.4 Qualitative Results
To intuitively illustrate the performance of X-Fi on the human sensing tasks, we visualize the estimated human skeleton generated by the model in the HPE task in Figure. 3 and employ the t-SNE method  (Van der Maaten & Hinton, 2008) to visualize the multimodal embedding obtained by the X-fusion block for the HAR task in Figure. 4.
@@
The visualization results for HPE comprise two actions, ‘picking up things’ and ‘throwing,’ each depicted through a sequence of four images. In different rows, we presented the results for the RGB and depth single-modal inputs, the fused results of the RGB and depth multimodal inputs, and the ground truth. To facilitate a clearer comparison between the fused results and the single-modal inputs results, we incorporated blue and orange dashed lines in the fused result images to represent the results of the RGB and depth single-modality inputs, respectively. Through observation, we found that the RGB modality provides more accurate estimations of leg movements, while the depth modality achieves greater accuracy in estimating hand movements. By combining both modalities into a multi-modal input, X-Fi effectively retains the leg movement accuracy from the RGB modality and the hand movement accuracy from the depth modality, resulting in fused outcomes that are closest to the ground truth.
@@
The visualization results for HAR includes the multi-modal embedding distributions from WiFi-CSI and mmWave Radar single-modal inputs, as well as the combined WiFi-CSI and mmWave Radar multi-modal inputs. To more closely analyze the distribution of sample points, we zoomed in on a small region containing points from two distinct categories. To quantify the distribution, we used the Silhouette score and Calinski–Harabasz index as indicators of clustering quality. The Silhouette score assesses how well an object fits within its own cluster compared to others, ranging from -1 to 1, with higher values indicating better separation and cohesion. Similarly, higher Calinski–Harabasz index values reflect more distinct clusters. We observed that the multimodal input results exhibit superior feature discriminability. While the mmWave Radar single-modal results demonstrate good classification capability, a closer inspection of the zoomed-in region reveals that the multimodal input results forms tighter clusters. Furthermore, a comparison of the indicator values confirms that the multimodal results outperforms any single-modal results.
@@ Appendix A Appendix
Table 5: Full results of the comparison among transformer-based fusion, X-Fusion, and X-Fusion’s variants.
A.1 Feature Extractor Pre-training.
As mentioned in the main work, to effectively extract informative high-dimensional representations from modal data, we designed modality-specific feature extractors whose backbone structures could be selected based on the data structure of each modality and the characteristics of the datasets. Meanwhile, we employ pre-training techniques to preserve the efficiency of feature extraction. Initially, for each modality, we train an entire model on the respective dataset and task. A portion of the model architecture is then repurposed as the feature extractor, with the corresponding pre-trained weights frozen during subsequent training. For instance, a ResNet-18 model was first trained on RGB images from the MM-Fi dataset to perform HPE. Subsequently, the MLP structure of the model’s output layer was discarded, retaining only the bottleneck blocks serving as the feature extractor. To identify efficient backbone structures, we conducted experiments to validate the performance of various backbone structures on each dataset.The finalized backbones are presented in the following. When utilizing MM-Fi for HPE, we employ the ResNet-18 architecture (He et al., 2016) for both the RGB and Depth modalities, the Point Transformer architecture (Zhao et al., 2021) for LiDAR, the Point Transformer architecture without transition down layers for mmWave point cloud data, and the MetaFi++ architecture (Zhou et al., 2023) for Wifi-CSI. For HAR using MM-Fi, the same backbone architectures are utilized, with the parameters specifically trained for HAR due to the differing feature representation requirements across tasks. When utilizing the XRF55 dataset for HAR, the ResNet-18 architecture is adopted for the RFID, WiFi, and mmWave modalities to maintain consistency with the baseline backbone networks employed in the study (Wang et al., 2024).

### v3

@@ 4.1 Experimental Setup
We train and evaluate our proposed X-Fi on the two largest human sensing multimodal public datasets, MM-Fi (Yang et al., 2024) and XRF55 (Wang et al., 2024), to assess its efficiency as a unified modality-invariant foundation model across diverse human sensing tasks, including Human Pose Estimation (HPE) and Human Activity Recognition (HAR).The dataset details for each task are as follows: (i) HPE: MM-Fi consists of over 320k synchronized frames from 40 human subjects. Each synchronized frame consists of samples collected from 5 sensing modalities, including RGB images (I), Depth images (D), LiDAR point clouds (L), mmWave point clouds (R), and WiFi-CSI (W). The chosen annotation type is 3D whole-body key points, representing the positions of 17 human joints. (ii) HAR: MM-Fi and XRF55 are utilized for HAR task evaluation. MM-Fi consists of 27 action categories, including 14 daily activities and 13 rehabilitation exercises. We exclude WiFi-CSI modality from MM-Fi when addressing the HAR task, as each WiFi data sample spans only 100ms, which is insufficient to effectively capture the distinct data patterns associated with different actions. XRF55 includes 42.9K radio frequency samples collected by 3 sensing modalities, including mmWave Range-Doppler & Range-Angle Heatmaps (R), WiFi-CSI (W), and RFID phase series data (RF). XRF55 comprises 55 action classes collected from 39 human subjects, covering categories such as fitness activities and human-computer interactions. We follow the S1 Random Split for MM-Fi and the original split setting for XRF55 as outlined in their respective papers.
@@
Different loss functions are applied as the training objectives for different human sensing tasks. For HPE, Mean-Squared-Error is chosen as the loss function, which measures the average squared euclidean distance between each of the 17 predicted whole-body key points and their corresponding ground truth points in 3D space. For HAR, Cross-Entropy loss is applied between the predicted probability distribution and the ground truth distribution.
Metrics MPJPE PA-MPJPE
@@
To evaluate the performance of our proposed unified modality-invariant foundation model across various human sensing tasks and modality combinations, we formulate baseline models by adopting a feature fusion approach (Das et al., 2021; Marcard et al., 2016) and a decision-level fusion approach (Yang et al., 2024). For the feature fusion approach, X-Fi uses pre-trained modality-specific extractors to obtain modality features. Single-modal inputs are projected into the task space via an MLP, while multi-modal inputs are concatenated and fused using an MLP with average pooling to map them into the unified task representation space. The baseline models are retrained on each modality combination separately. For different tasks, we retrain baseline models’ weights (including the pre-trained feature extractor) using different learning objectives. For the decision-level fusion approach, we first obtain the final-layer output of each individual modality. Then, we calculate the average of the outputs from all modalities in the combination to generate the fused output. The baseline results are presented in the Table 12.
@@
We conduct experiments on all human sensing tasks and datasets with the following model settings. To standardize feature representations obtained by various modality-specific feature extractors, we apply linear projection units to map each modality feature representation to $n_{f}=32$, each with a feature dimension of $d_{f}=512$. The positional embedding, derived from LiDAR 3D point cloud raw data due to its inherent capture of human body spatial information, is only provided in the multi-modal embedding if the LiDAR modality is present in the input. The backbone for both the cross-modal transformer and each modality-specific cross-attention module consists of a 1-layer decoder-only transformer structure with 8 multi-head attention heads and a scaling factor of 0.125. Each modality-specific key-value pair and the cross-modal embedding in X-Fusion block comprise 32 features with a feature dimension of 512. The number of iterations on the X-Fusion block is set to a default of 4 in our experiments. The AdamW optimizer, with an initial learning rate of $1\times e^{-3}$ for HPE and $1\times e^{-4}$ for HAR, is chosen for model optimization. The training process is performed with a batch size of 16 on an NVIDIA GeForce RTX 4090 GPU.
@@ 4.2 Quantitative Results
Table 2: HAR accuracy (%) on the MM-Fi and XRF55 datasets.
@@
R 82.1 83.9 1.9%
W 77.8 55.7 -22.1%
RF 42.2 42.5 0.3%
R+W 86.8 88.2 1.4%
R+RF 71.4 86.5 15.1%
W+RF 55.6 58.1 2.6%
R+W+RF 70.6 89.8 19.2%
We compare our approach to baseline methods in downstream human sensing tasks, including HPE and HAR. X-Fi results are obtained by evaluating a single X-Fi model, only trained once using the proposed modality-invariant training strategy, across various modality combinations during testing. And the baseline results are retrained on each modality combination separately. For convenience, we refer to the decision-level fusion baseline as “baseline1”, and the feature-level fusion baseline as “baseline2”. The results for X-Fi, as shown in Tables 12, represent average values obtained from multiple experiments.
Human Pose Estimation.
We assess the performance of X-Fi on the MM-Fi dataset (Yang et al., 2024) by adopting two commonly-used metrics in HPE: Mean Per Joint Position Error (MPJPE) and Procrustes Analysis MPJPE (PA-MPJPE). For single-modal inputs, X-Fi improves the averaged MPJPE and PA-MPJPE by $6.0\%$ and $5.2\%$, respectively. For the dual-modality inputs, the averaged MPJPE and PA-MPJPE are improved respectively by $14.5\%$ and $14.8\%$ on baseline1, $23.7\%$ and $18.9\%$ on baseline2. Notably, for LiDAR and WiFi modalities, when used as individual inputs, X-Fi does not achieve good performance, with MPJPE difference of $-3.5\%$ and $+0.7\%$, respectively. However, when the two modalities are combined, the performance improves significantly, surpassing the baseline1 by $4.5\%$ and the baseline2 by $22.6\%$. For the triple-modality inputs, X-Fi demonstrates an average improvement of $17.0\%$ and $28.8\%$ in MPJPE, $19.5\%$ and $25.9\%$ in PA-MPJPE. For the combined input of four modalities, the results indicate an average enhancement of $17.0\%$ and $35.5\%$ in MPJPE, $21.6\%$ and $30.8\%$ in PA-MPJPE. For all five modalities combined input, MPJPE increases by $18.8\%$ and $35.1\%$, PA-MPJPE increases by $23.8\%$ and $34.3\%$.
@@
The performance of X-Fi on the HAR task is evaluated by classification accuracy on both the MM-Fi (Yang et al., 2024) and XRF55 (Wang et al., 2024) datasets. On the MM-Fi dataset, X-Fi performs better than the baseline when handling multimodal inputs, yielding an average improvement of 2.8%. But when dealing with single-modality inputs, its performance declines, dropping by 2.8%. On the XRF55 dataset, X-Fi only encounters challenges with WiFi-CSI single-modality input. It outperforms the baseline on all other modality combinations, achieving an average improvement of 6.7%. Notably, while X-Fi’s performance is suboptimal when using WiFi as a single input, when combined with other modalities, WiFi not only avoids hindering performance but also introduces unique information that enhances classification accuracy. For example, when X-Fi processes mmWave radar and RFID multimodal input, it outperforms the baseline by 15.1%. With the inclusion of WiFi, this improvement increases to 19.2%.
@@ 4.3 Analytics
To assess the impact of adjusting the independent existing probability of each modality in the modality existence list, we conduct experiments with different probability combinations on the XRF55 dataset for the HAR task. The XRF55 modality existence list contains the independent existence probabilities of mmWave Radar, WiFi-CSI, and RFID modalities in sequential order. As shown in Table 3, X-Fi encounters difficulties in dealing with WiFi-CSI. However, through gradually increasing the existing probability of WiFi-CSI, the model is prioritized to fit WiFi-CSI modality during training, which yields significant improvement from $29.1\%$ to $55.7\%$ in handling single WiFi-CSI input. Additionally, we are surprised to find that prioritizing a particular modality does not degrade the model’s ability to extract features from the other two modalities. Consequently, the multimodal fusion results are also improved a lot.
@@
Method VO Transformer X-Fi VO Transformer X-Fi
@@
Table 4: Comparison between the SOTA method (Visual Odometry Transformer) and X-Fi on the HPE task.
Comparison between existing SOTA (Visual Odometry Transformer) and X-Fi.
To demonstrate the efficiency of X-Fi on modality-invariant human sensing tasks, we conduct comparison with existing SOTA method (Memmel et al., 2023) that utilize basic transformer encoder consisting of four stacked blocks to performance modality fusion. A subset of comparison results are analyzed in Table 4 and the full findings are provided in the appendix 5. As shown in Table 4, X-Fusion yields better performance on most of the modality combination input. First, when handling single modality input, transformer-based fusion struggles with challenging modalities, whereas X-Fusion effectively captures the distinctive features of each modality. For example, performance with LiDAR input improves by $21.3\%$. Next, in contrast to transformer-based fusion methods, X-Fusion leverages each modality’s strengths, yielding better fusion results than single-modality inputs. X-Fusion demonstrates superior performance compared to either the LiDAR or WiFi-CSI modality when used individually. Lastly, in the case of multi-modality inputs, the performance of transformer-based fusion is highly dependent on the presence of dominant modalities. For example, adding the RGB modality improves the performance of the LiDAR and WiFi-CSI multi-modal inputs by 119.7 mm with the transformer backbone, whereas the corresponding improvement with X-Fusion is 66.6 mm.
@@
To assess the effectiveness of the X-Fusion structure, which iterates multiple loops within the same block for multimodal feature fusion, we design two variants. The first employs a layered architecture with four sequentially stacked X-Fusion blocks, where each layer learns new key-value pairs from the multimodal embedding of the previous layer. The second variant is similar but uses the same key-value pairs across layers, learned from the embedding generated by the modality feature encoder. We conduct experiments on both variants, with full results provided in the appendix 5. On average, the first variant demonstrates a performance improvement of +1.27 mm for MPJPE and -0.7 mm for PA-MPJPE across all modality combinations. The second variant achieves an improvement of 0.93 mm for MPJPE and 0.03 mm for PA-MPJPE. Considering the computational efficiency and the minor performance differences, we retain the X-Fusion structure where an X-Fusion block is applied iteratively.
@@ 4.4 Qualitative Results
To intuitively illustrate the performance of X-Fi on the human sensing tasks, we visualize the estimated human skeleton generated by the model in the HPE task in the Figure 3 and employ the t-SNE method  (Van der Maaten & Hinton, 2008) to visualize the multimodal embedding obtained by the X-fusion block for the HAR task in the Figure 4.
@@
The visualization results for HPE comprise two actions, ‘picking up things’ and ‘throwing’, each depicted through a sequence of four images. In different rows, we presented the results for the RGB and depth single-modal inputs, the fused results of the RGB and depth multimodal inputs, and the ground truth. To facilitate a clearer comparison between the fused results and the single-modal inputs results, we incorporated blue and orange dashed lines in the fused result images to represent the results of the RGB and depth single-modality inputs, respectively. Through observation, we found that the RGB modality provides more accurate estimations of leg movements, while the depth modality achieves greater accuracy in estimating hand movements. By combining both modalities into a multimodal input, X-Fi effectively retains the leg movement accuracy from the RGB modality and the hand movement accuracy from the depth modality, resulting in fused outcomes that are closest to the ground truth.
@@
The visualization results for HAR include the multimodal embedding distributions from WiFi-CSI and mmWave Radar single-modal inputs, as well as the combined WiFi-CSI and mmWave Radar multi-modal inputs. To more closely analyze the distribution of sample points, we zoomed in on a small region containing points from two distinct categories. To quantify the distribution, we used the Silhouette score and Calinski–Harabasz index as indicators of clustering quality. The Silhouette score assesses how well an object fits within its own cluster compared to others, ranging from -1 to 1, with higher values indicating better separation and cohesion. Similarly, higher Calinski–Harabasz index values reflect more distinct clusters. We observed that the multimodal input results exhibit superior feature discriminability. While the mmWave Radar single-modal results demonstrate good classification capability, a closer inspection of the zoomed-in region reveals that the multimodal input results form tighter clusters. Furthermore, a comparison of the indicator values confirms that the multimodal results outperform any single-modal results.
@@ Methods
Table 5: Full results of the comparisons among Visual-Odometry Transformer based fusion , X-Fusion, and X-Fusion’s variants.
A.1 Feature Extractor Pre-training
As mentioned in the main work, to effectively extract informative high-dimensional representations from modal data, we designed modality-specific feature extractors whose backbone structures could be selected based on the data structure of each modality and the characteristics of the datasets. Meanwhile, we employ pre-training techniques to preserve the efficiency of feature extraction. Initially, for each modality, we train an entire model on the respective dataset and task. A portion of the model architecture is then repurposed as the feature extractor, with the corresponding pre-trained weights frozen during subsequent training. For instance, a ResNet-18 model was first trained on RGB images from the MM-Fi dataset to perform HPE. Subsequently, the MLP structure of the model’s output layer was discarded, retaining only the bottleneck blocks serving as the feature extractor. To identify efficient backbone structures, we conducted experiments to validate the performance of various backbone structures on each dataset. The finalized backbones are presented below. When utilizing MM-Fi for HPE, we employ the ResNet-18 architecture (He et al., 2016) for both the RGB and Depth modalities, the Point Transformer architecture (Zhao et al., 2021) for LiDAR, the Point Transformer architecture without transition down layers for mmWave point cloud data, and the MetaFi++ architecture (Zhou et al., 2023) for Wifi-CSI. For HAR using MM-Fi, the same backbone architectures are utilized, with the parameters specifically trained for HAR due to the differing feature representation requirements across tasks. When utilizing the XRF55 dataset for HAR, the ResNet-18 architecture is adopted for the RFID, WiFi, and mmWave modalities to maintain consistency with the baseline backbone networks employed in the study (Wang et al., 2024).